### v2

Abstract
Patch-based transformers have emerged as efficient and improved long-horizon modeling architectures for time series modeling. Yet, existing approaches rely on temporally-agnostic patch construction, where arbitrary starting positions and fixed lengths fracture temporal coherence by splitting natural transitions across boundaries. This naive segmentation often disrupts short-term dependencies and weakens representation learning. We propose a novel Entropy-Guided Dynamic Patch Encoder (EntroPE), as a temporally informed framework that dynamically detects transition points via conditional entropy and dynamically places patch boundaries. This preserves temporal structure while retaining the computational benefits of patching. EntroPE consists of two key modules, namely an Entropy-based Dynamic Patcher (EDP) that applies information-theoretic criteria to locate natural temporal shifts and determine patch boundaries, and an Adaptive Patch Encoder (APE) that employs pooling and cross-attention to capture intra-patch dependencies and produce fixed-size latent representations. Extensive experiments on long-term forecasting, classification, and anomaly detection demonstrate that the proposed method improves both accuracy and efficiency, establishing entropy-guided dynamic patching as a promising new paradigm for time series modeling.
1 Introduction
Figure 1: Comparison of patching strategies. (a) Temporally-agnostic fixed-length patching places boundaries at arbitrary positions, fragmenting coherent temporal patterns (highlighted regions show split transitions). (b) Our temporally-informed approach places boundaries at entropy peaks, where conditional entropy scale indicates natural transition points, preserving temporal coherence within patches.
Time series analysis is fundamental to numerous critical applications, including electricity load forecasting (Gasparin et al., 2022), financial market prediction (Fischer and Krauss, 2018), healthcare monitoring (Reis and Mandl, 2003), and environmental surveillance (Smith, 1989). However, accurately modeling time series data remains challenging due to inherent noise, complex temporal dependencies, and irregular patterns spanning multiple time horizons (Lim and Zohren, 2021; Torres et al., 2021). The transformer architecture (Vaswani et al., 2017) has revolutionized language and sequence modeling across domains through self-attention mechanisms, leading to significant advances in time series analysis. Despite their promise in learning long-term temporal relationships, time series transformers remain suboptimal due to unique complexities such as non-stationarity, seasonality, varying temporal granularities (Wen et al., 2023; Zeng et al., 2023).
Recent advances in time series transformers include Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022), which address long-sequence modeling challenges through sparse attention mechanisms, decomposition techniques, and improved positional encodings while operating on point-wise inputs. Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) introduced a patch-based input representation, dividing raw sequences into fixed-length patches and achieving substantial improvements in computational efficiency and performance. This demonstrated the critical impact of input tokenization strategies on transformer performance in time series. However, subsequent approaches have predominantly adopted similar temporally-agnostic patching schemes, which may not fully capture the temporal coherence and statistical properties inherent in time series data.
These patching strategies treat time as temporally-homogeneous, partitioning sequences at arbitrary positions without regard for underlying temporal structure. Temporally-agnostic patching introduces two fundamental challenges. They create inconsistent input representations between training and inference phases. While training observes patches starting at diverse temporal offsets, result in patches that begin at various temporal positions relative to underlying patterns (e.g., a patch might start mid-trend in one sample but at a trend beginning in another). However, during inference, patches are extracted from predetermined positions in the input sequence, creating a systematic distribution shift in patch configurations. Further, predetermined patch boundaries arbitrarily segment coherent temporal structures without regard for frequent underlying dynamics (see Fig. 1(a)). For example, a frequently occurring gradual trend change or seasonal transition may be split across multiple patches, breaking the natural temporal dependencies that are crucial for accurate pattern recognition.
These challenges directly impact model performance. Fragmented temporal patterns lead to incomplete intra-patch representations, where semantically related time points are separated across different patches and processed independently. Additionally, the train-inference mismatch reduces the model’s ability to generalize, as it must extrapolate to patch configurations with different statistical properties than those encountered during training. This is particularly detrimental for capturing rapid transitions and fine-grained temporal dynamics that require consistent temporal context.
We argue that patch boundaries should align with points of elevated predictive uncertainty, where the past provides insufficient information to determine the future. Formally, such moments are characterized by high conditional entropy of the next observation given historical context. Segmenting time at these entropy peaks localizes regime changes at patch boundaries while yielding patches whose interiors remain predictively coherent. Motivated by this principle, we introduce EntroPE, an Entropy-Guided Dynamic Patch Encoder, which integrates uncertainty-aware boundary detection as a core architectural component. Unlike prior approaches that fix patch length or rely on heuristic change-point detectors, EntroPE derives segmentation from information-theoretic predictive structure, ensuring patch boundaries align with natural temporal transitions (Fig. 1(b)). We then employ adaptive encoding mechanisms to process these variable-length patches while preserving intra-patch dependencies. This temporally-informed approach addresses both the train-inference mismatch and boundary fragmentation issues inherent in temporally-agnostic patching methods.
The main contributions of this paper are:
  • We propose EntroPE, the first information-theoretic framework for dynamic patching in time series transformers.
  • We introduce an adaptive segmentation mechanism that balances modeling accuracy and attention complexity by dynamically controlling patch resolution.
  • We demonstrate consistent improvements across forecasting, classification, and anomaly detection, establishing entropy-guided dynamic patching as a practical and generalizable paradigm for time series analysis.
Code is publicly available at https://github.com/Sachithx/EntroPE.
2 Related Work
Transformers have significantly advanced time series modeling by enabling long-range dependency learning through self-attention. Early adaptations such as Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022) improved scalability via sparse attention, decomposition, and frequency-domain modeling. However, these approaches operate on point-wise tokens, resulting in prohibitively long sequences and a fundamental trade-off between temporal resolution and computational tractability.
Patch-Based Input Representations. Inspired by Vision Transformers (Dosovitskiy et al., 2021), patch-based methods reduce sequence length while preserving local temporal semantics. Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) pioneered fixed-length temporal patching, demonstrating substantial gains in efficiency and forecasting accuracy. Subsequent work extended this paradigm through hierarchical dependency modeling (Crossformer (Zhang and Yan, 2023)), multivariate forecasting adaptation (CARD (Xue et al., 2024)), and patch-independent self-supervised learning (Lee et al., 2024). Beyond transformers, xPatch (Stitsyuk and Choi, 2025), Pathformer (Chen et al., 2024), and PatchMLP (Tang and Zhang, 2025) explore CNN- and MLP-based hierarchical architectures to capture multi-scale temporal patterns.
Adaptive and Multi-Granularity Patching. Recent efforts relax fixed patch sizes through adaptive or multi-scale segmentation. MSPatch (Cao et al., 2025) employs multi-resolution patches, HDMixer (Huang et al., 2024) enables length-extendable interpolation, and MOIRAI (Woo et al., 2024) determines patch size in the frequency domain. AdaPatch (Liu et al., 2025a) mitigates non-stationarity via patch-level normalization and reconstruction, but operates over fixed segmentation rather than learning boundary placement. APN (Liu et al., 2025c) introduces time-aware adaptive patching for irregular time series through soft window aggregation, emphasizing data-density regularization instead of causality driven segmentation. SRSNet (Wu et al., 2025) constructs a selective representation space by scoring, selecting, and reassembling candidate patches, but operates over dense sliding-window patch pools rather than discovering causally meaningful boundary locations. TimeMosaic (Ding et al., 2025) further adapts patch granularity based on temporal heterogeneity and motif reuse, optimizing representation efficiency and decoding specialization through segment-wise prompting. While these methods improve flexibility, their segmentation criteria remain largely heuristic or representation-driven, prioritizing encoding density, frequency structure, or compression objectives over the underlying predictive or causal structure of the time series.
Causal and Information-Theoretic Perspectives on Segmentation. A complementary line of research argues that segmentation should reflect changes in predictive structure. In NLP, Byte Latent Transformer (Pagnoni et al., 2025) demonstrates that architecture-integrated dynamic tokenization improves efficiency by adapting token boundaries to information content. However, existing time series methods have not systematically grounded patch boundary selection in information-theoretic measures of predictive uncertainty or temporal causality.
In contrast to prior work that optimizes patching for representational efficiency or frequency structure, EntroPE places patch boundaries at peaks of conditional predictive entropy. This identifies moments where the past becomes insufficient to predict the future, typically corresponding to regime shifts or structural transitions. This principled segmentation preserves temporal coherence within patches, localizes uncertainty at boundaries, and eliminates train-inference patch mismatch. By coupling entropy-guided dynamic patching with adaptive encoding for variable-length segments, our method introduces a fundamentally new criterion for patch construction based on predictive uncertainty rather than heuristic granularity.
3 Method
3.1 Preliminaries: Problem Formulation
Given a multivariate time series $X=[x^{(1)},\ldots,x^{(C)}]\in\mathbb{R}^{C\times L}$ with $C$ channels and look-back length $L$, we address three fundamental tasks, namely forecasting, classification, and anomaly detection. Following the channel-independence principle (Nie et al., 2023), our architecture $f_{\phi}$ processes each channel independently to obtain channel-wise representations $\mathbf{z}^{(c)}=f_{\phi}(x^{(c)})$ for $c=1,\ldots,C$. Here, $f_{\phi}$ encompasses the complete processing pipeline: entropy-guided dynamic patching, adaptive patch encoding, global transformer, and fusion decoder (detailed in Sec. 3.3-3.6), producing representations $Z=[\mathbf{z}^{(1)},\ldots,\mathbf{z}^{(C)}]\in\mathbb{R}^{C\times d}$ where $d$ is the final representation dimension. $Z$ is projected via linear layer based on specific task.
Figure 2: Comprehensive architecture of EntroPE. The model processes input through: (A) Entropy-Based Dynamic Patcher - A small causal transformer calculates entropy at each time point to identify boundaries where predictive uncertainty is high; (B) Adaptive Patch Encoder - Cross-attention layers aggregate intra-patch dependencies into fixed-size global embeddings; (C) Fusion Decoder - Cross-attention combines global patch context with local encoder hidden states for accurate predictions.
3.2 Overall Architecture
Our approach consists of two distinct phases (Fig. 2), (1) entropy model pre-training for uncertainty quantification, (2) task-specific training with entropy guided dynamic patching. Crucially, the entropy model parameters remain frozen during task-specific training, serving solely to provide patch boundaries.
3.3 Entropy Model Pre-training
As shown in Fig. 2(A), our dynamic patch boundary detection leverages a lightweight causal transformer to identify temporal transitions based on predictive uncertainty. Given a time series dataset $\mathcal{D}_{\text{pretrain}}=\{X^{(i)}\}_{i=1}^{N}$ where $X^{(i)}\in\mathbb{R}^{C\times L}$, we first quantize each continuous sequence into discrete tokens. For a univariate sequence $x=[x_{1},\ldots,x_{L}]$, we apply quantization $q:\mathbb{R}\to\mathcal{V}$ (defined in Eq. 38) to obtain tokenized sequence $\tau=[\tau_{1},\ldots,\tau_{L}]$ where $\tau_{t}=q(x_{t})\in\mathcal{V}$ and $|\mathcal{V}|=V$.
A critical distinction. Unlike Chronos (Ansari et al., 2024), which operates entirely in the quantized domain and suffers from range limitations during dequantization, we use quantization only for computing entropies and input projection. The downstream prediction heads project the output-embeddings, not quantized tokens. This decoupling eliminates the quantization range constraint as our predictions are not limited to the discrete vocabulary. Quantization serves purely as a preprocessing step for the entropy model (Sec. 3.3), while the actual time series transformer and prediction heads operate on continuous representations throughout.
Determining Quantization Range. To avoid manual threshold selection, we determine bin boundaries $\{b_{1},\ldots,b_{V-1}\}$ from training data statistics. Let $X_{\text{train}}$ denote z-score normalized training values. We compute empirical quantiles:
$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\quad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}})$ (1)
where $\epsilon=0.005$ captures 99.5% of the distribution. The symmetric quantization radius is $R=\max(|q_{\text{low}}|,|q_{\text{high}}|)$, and the interval [R,R][-R,R] is uniformly divided into $V$ bins.
Model Selection. Drawing inspiration from decoder-only transformers in time series (Das et al., 2024) and cross-entropy guided training in Chronos, we adopt the GPT-2 paradigm (Radford et al., 2019) to train a compact transformer for next-token prediction on quantized time series tokens. Our lightweight architecture comprises 8-16 embedding dimensions and 2 transformer layers, totaling $\approx 10$k learnable parameters. The model performs autoregressive prediction at timesteps $\{t_{2},t_{3},\ldots,t_{L}\}$ from input sequence $\{t_{1},t_{2},\ldots,t_{L-1}\}$.
Model Training. We train our lightweight causal transformer $f_{\theta_{\text{ent}}}:\mathcal{V}^{t}\to\Delta^{V}$ (where $\Delta^{V}$ denotes the probability simplex over vocabulary $\mathcal{V}$) to perform next-token prediction. The model is optimized via:
$\theta_{\text{ent}}^{*}=\arg\min_{\theta_{\text{ent}}}\mathbb{E}_{X\sim\mathcal{D}_{\text{pretrain}}}\left[\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})\right]$ (2)
where $\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})$ is the cross-entropy loss (Eq. 21).
Once converged, we freeze $\theta_{\text{ent}}^{*}$ and use it exclusively for entropy computation, not for prediction or token generation. This decoupling is critical, as the entropy model learns predictive uncertainty patterns, while downstream task models operate on continuous representations.
3.4 Entropy-Guided Dynamic Patcher (EDP)
Entropy Computation.
For a new input sequence $X\in\mathbb{R}^{C\times L}$, we process each channel independently. For channel $c$, we compute the quantized sequence $\tau^{(c)}=[q(x_{1}^{(c)}),\ldots,q(x_{L}^{(c)})]$ and use the frozen model $\theta_{\text{ent}}^{*}$ to calculate conditional entropy at each position:
$H(x_{t}^{(c)};\theta_{\text{ent}}^{*})=\\ -\sum_{v\in\mathcal{V}}p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})\log p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})$ (3)
yielding entropy sequence,
$\mathbf{H}^{(c)}=[H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})]\in\mathbb{R}^{L}.$ (4)
As established in Appendix A.2, high conditional entropy indicates positions where the future is less predictable from the past, typically at regime shifts, trend reversals, or volatility bursts. By placing boundaries at such points (Proposition A.2), patches become internally coherent while transition uncertainty is localized at edges.
Adaptive Threshold Determination.
Rather than using fixed global thresholds requiring manual tuning, we employ sample-adaptive thresholds derived from quantile-based estimation. For each input sequence with entropies $\{H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})\}$, we compute sample-specific thresholds. Hence, the threshold selection becomes interpretable.
The absolute entropy threshold is the $\alpha$-th quantile of the sequence’s entropy distribution:
$\theta_{\text{abs}}^{(c)}=Q_{\alpha}(\mathbf{H}^{(c)}).$ (5)
The relative threshold is the $\alpha$-th quantile of consecutive differences $\Delta\mathbf{H}^{(c)}=\bigl[H(x_{t}^{(c)})-H(x_{t-1}^{(c)})\bigr]_{t=2}^{L}$,
$\gamma_{\mathrm{rel}}^{(c)}$ $=Q_{\alpha}\!\left(\Delta\mathbf{H}^{(c)}\right),$ (6)
where $\alpha$ is the quantile parameter (e.g., $\alpha=75\%$). This formulation enables automatic adaptation to varying entropy levels across sequences, eliminating dataset-specific threshold tuning.
Boundary Detection.
A patch boundary is placed at position $t$ based on the following conditions:
$H(x_{t}^{(c)})$ $>\theta_{\text{abs}}^{(c)}\quad\text{(high absolute uncertainty)}$ (7)
$\Delta H(x_{t}^{(c)})$ $>\gamma_{\text{rel}}^{(c)}\quad\text{(significant uncertainty increase)}$ (8)
$M_{t-1}^{(c)}$ $=0\quad\text{(no consecutive boundaries)}.$ (9)
This threshold mechanism produces a binary patch boundary mask $M^{(c)}\in\{0,1\}^{L}$ where $M_{t}^{(c)}=1$ indicates a patch start. The framework supports using either threshold independently or in combination, allowing flexible adaptation to different temporal characteristics. EDP outputs mask $M$ for subsequent non-overlapping patch processing.
Patch Segmentation.
Given boundary mask $M^{(c)}$, we extract boundary positions $\mathcal{B}^{(c)}=\{t:M_{t}^{(c)}=1\}=\{b_{1}^{(c)},b_{2}^{(c)},\ldots,b_{P^{(c)}+1}^{(c)}\}$ where $b_{1}^{(c)}=1$ and $b_{P^{(c)}+1}^{(c)}=L+1$. The sequence is then partitioned into $P^{(c)}$ non-overlapping patches:
$x^{(c)}=\bigcup_{j=1}^{P^{(c)}}p_{j}^{(c)},$ (10)
where $\quad p_{j}^{(c)}=[x_{b_{j}^{(c)}},x_{b_{j}^{(c)}+1},\ldots,x_{b_{j+1}^{(c)}-1}]\in\mathbb{R}^{\ell_{j}^{(c)}}$, $\ell_{j}^{(c)}=b_{j+1}^{(c)}-b_{j}^{(c)}$ is the length of patch $j$, and $\sum_{j=1}^{P^{(c)}}\ell_{j}^{(c)}=L$. Our segmentation depends only on relative uncertainty structure, not absolute probability calibration.
3.5 Adaptive Patch Encoder (APE)
The EDP produces variable-length patches where both patch count and individual lengths vary across samples. As shown in Fig. 2(B), the APE converts these heterogeneous patches into fixed-size representations suitable for transformer processing while preserving temporal information. When handling variable patches, we ensure batch processing compatibility by padding sequences to the maximum patch count within each batch and applying attention masks to handle variable patch lengths during cross-attention operations.
Architecture Design. Our encoder employs a two-stage approach. First initial dimensionality reduction via pooling (from time-point embeddings to initial patch embeddings), followed by iterative cross-attention refinement. This design choice is inspired by the Perceiver architecture (Jaegle et al., 2021) and Byte Latent Transformer (Pagnoni et al., 2025), due to their demonstrated performance on language and image data.
Patch Encoding.
Each variable-length patch $p_{j}^{(c)}$ is encoded into a fixed-dimensional representation via the APE. First, we obtain time-point embeddings through an embedding layer $E:\mathcal{V}\to\mathbb{R}^{d_{t}}$:
$\mathbf{h}_{t}^{(c)}=E(q(x_{t}^{(c)})),\quad t=1,\ldots,L.$ (11)
For patch $p_{j}^{(c)}$ spanning positions $\mathcal{T}_{j}^{(c)}=\{b_{j}^{(c)},b_{j}^{(c)}+1,\ldots,b_{j+1}^{(c)}-1\}$, we initialize its representation via max pooling:
$\mathbf{z}_{j}^{(c),0}=\text{MaxPool}\left(\{\mathbf{h}_{t}^{(c)}\}_{t\in\mathcal{T}_{j}^{(c)}}\right)\in\mathbb{R}^{d_{p}}.$ (12)
We then refine patch embeddings through $N$ cross-attention layers where patch embeddings query only their constituent time points:
$\mathbf{z}_{j}^{(c),n}=\mathbf{z}_{j}^{(c),n-1}+W_{o}\sum_{i\in\mathcal{T}_{j}^{(c)}}\alpha_{ji}^{(n)}W_{v}(\mathbf{h}_{i}^{(c),n-1})$ (13)
where attention weights are computed only within the patch:
$\alpha_{ji}^{(n)}=\frac{\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i}^{(c),n-1})}{\sqrt{d_{k}}}\right)}{\sum_{i^{\prime}\in\mathcal{T}_{j}^{(c)}}\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i^{\prime}}^{(c),n-1})}{\sqrt{d_{k}}}\right)}$ (14)
This produces final patch embeddings $P^{(c)}=[\mathbf{z}_{1}^{(c),N},\ldots,\mathbf{z}_{P^{(c)}}^{(c),N}]\in\mathbb{R}^{P^{(c)}\times d_{p}}$ and encoder hidden states $H^{(c)}=[\mathbf{h}_{1}^{(c),N},\ldots,\mathbf{h}_{L}^{(c),N}]\in\mathbb{R}^{L\times d_{t}}$.
Batch Processing with Variable Patches.
For a batch of $B$ samples with varying patch counts $\{P^{(b,c)}\}$, let $P_{\max}=\max_{b,c}P^{(b,c)}$. We construct a padded patch embedding tensor $\tilde{P}^{in}\in\mathbb{R}^{(B\cdot C)\times P_{\max}\times d_{p}}$ with attention mask $\mathcal{M}\in\{0,1\}^{(B\cdot C)\times P_{\max}}$ where $\mathcal{M}_{bc,j}=\mathbb{1}[j\leq P^{(b,c)}]$ to mask padded positions during self-attention.
3.6 Global Transformer & Fusion Decoder
The final components of the EntroPE process are the fixed-size patch embeddings to learn long-term dependencies and generate predictions.
Global Transformer.
Patch embeddings are processed through a standard transformer to capture inter-patch dependencies:
$\tilde{P}^{(c)}=\text{Transformer}(P^{(c)},\mathcal{M})\in\mathbb{R}^{P^{(c)}\times d_{p}}$ (15)
Fusion Decoder.
We fuse global patch context with fine-grained temporal information via cross-attention where time-point embeddings query patch representations:
$\tilde{H}^{(c)}=H^{(c)}+\\ W_{o}\left(\text{softmax}\left(\frac{[W_{q}(H^{(c)})][W_{k}(\tilde{P}^{(c)})]^{\top}}{\sqrt{d_{k}}}\right)W_{v}(\tilde{P}^{(c)})\right)\\ \in\mathbb{R}^{L\times d_{t}}$ (16)
This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, preserving both local temporal patterns and global contextual dependencies.
Task-Specific Head. The enriched representations $\tilde{H}^{(c)}$ are projected to generate future predictions. We apply flattening followed by a linear projection with Instance De-Normalization: $\text{InstanceDeNorm}\!\left(W_{\text{proj}}(\text{Flatten}(\tilde{H}^{(c)}))\right)$. Training criteria and additional details on task-specific heads are provided in Appendix B.
4 Experiments
In the section we evaluate the efficacy of EntroPE on long-term forecasting, classification and anomaly detection tasks. We show that our model can serve as a foundation model with competitive performance accross these tasks.
4.1 Long-Term Forecasting
Datasets. To examine forecasting performance we conduct comprehensive evaluations on six widely-used long-term multivariate forecasting datasets. i.e., Electricity (ECL) featuring electricity consumption data, ETT family (ETTh1, ETTh2, ETTm1, ETTm2) that encompass a range of scenarios in energy transfer technology, Weather that offers insights into various meteorological variables over time. Detailed descriptions in Appendix Tab. 6.
Baselines and Experimental Settings. We benchmark against fifteen state-of-the-art baselines. (1) Transformer-based models: TimeMosaic (Ding et al., 2025), iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021); (2) CNN-based model: TimesNet (Wu et al., 2023), TSLANet (Eldele et al., 2024); (3) MLP-based models: TimeMixer (Wang et al., 2024), HDMixer (Huang et al., 2024), DLinear (Zeng et al., 2023); (4) Foundation and lightweight models: Time-FFM (Liu et al., 2024a), TimeBase (Huang et al., 2025a); (5) LLM-aligned and cross-modal models: LangTime (Niu et al., 2025), CALF (Liu et al., 2025b); (6) Filter-based model: FilterTS (Wang et al., 2025). (Further details on baselines are discussed in Appendix D.1.)
Table 1: Multivariate time series forecasting results on benchmark datasets. Results are averaged across prediction horizons $T=\{96,192,336,720\}$ with fixed input length $L=96$ for all datasets. Best results are highlighted in red and second/-best in blue.
Models ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Autoformer [2021] 0.496 0.487 0.450 0.459 0.588 0.517 0.327 0.371 0.338 0.382 0.227 0.338
FEDformer [2022] 0.498 0.484 0.437 0.449 0.448 0.452 0.305 0.349 0.309 0.360 0.214 0.327
DLinear [2023] 0.461 0.457 0.563 0.519 0.404 0.408 0.354 0.402 0.265 0.315 0.225 0.319
TimesNet [2023] 0.495 0.450 0.414 0.427 0.400 0.406 0.291 0.333 0.251 0.294 0.193 0.304
PatchTST [2023] 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318
Time-FFM [2024] 0.442 0.434 0.382 0.406 0.399 0.402 0.286 0.332 0.270 0.288 0.216 0.299
HDMixer [2024] 0.448 0.437 0.384 0.407 0.396 0.402 0.286 0.331 0.253 0.285 0.205 0.295
iTransformer [2024] 0.454 0.447 0.383 0.407 0.407 0.410 0.288 0.332 0.258 0.278 0.178 0.270
TimeMixer [2024] 0.459 0.444 0.390 0.409 0.382 0.397 0.279 0.324 0.245 0.276 0.182 0.272
TSLANet [2024] 0.448 0.441 0.372 0.399 0.378 0.397 0.283 0.327 0.259 0.280 0.199 0.283
TimeBase [2025] 0.463 0.429 0.409 0.425 0.431 0.420 0.290 0.332 0.252 0.279 0.227 0.296
LangTime [2025] 0.437 0.425 0.375 0.392 0.397 0.392 0.284 0.321 0.252 0.273 0.201 0.285
FilterTS [2025] 0.440 0.432 0.375 0.399 0.386 0.397 0.279 0.323 0.253 0.280 0.184 0.275
CALF [2025] 0.441 0.435 0.372 0.395 0.396 0.391 0.280 0.321 0.250 0.274 0.177 0.266
TimeMosaic [2026] 0.427 0.442 0.369 0.399 0.388 0.392 0.278 0.319 0.253 0.277 0.188 0.281
EntroPE [2026] 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271
Results.
Table 1 shows EntroPE achieves competitive or superior performance across all baseline categories (Transformer, CNN, MLP, foundation, LLM-aligned, and filter-based models). While retaining PatchTST-like architecture, EntroPE introduces dynamic patching and adaptive encoding, yielding  20% MSE reduction on ETTh1,  15% on Electricity, and  10% average improvement over PatchTST, alongside reduced token count through non-overlapping patches. Notably, iTransformer and CALF demonstrates remarkable performance on high-dimensional Electricity dataset (321 variables) due to its channel-wise attention design. While CALF ($\sim$18M parameters) and LangTime ($\sim$500M parameters) show competitive results, EntroPE achieves comparable accuracy with 500-1000× fewer parameters (100k-1M), demonstrating that entropy-guided segmentation provides robust improvements without requiring massive model capacity.
Compared to some noteworthy dynamic or adaptive patching approaches such as HDMixer and TimeMosaic, EntroPE consistently delivers stronger or more stable gains across datasets. Notably, TimeMosaic employs substantially larger models (3–13M parameters) and higher training durations, whereas EntroPE achieves comparable or better accuracy with smaller models. This highlights that performance gains arise from principled entropy-driven boundary placement rather than increased model capacity. (Full table: Tab. 16 & 17)
4.2 Classification
Datasets. To demonstrate the generalizability of EntroPE beyond forecasting, we evaluate on 10 UEA multivariate time series classification datasets (Bagnall et al., 2018). We replace the forecasting head with a classification head while keeping the core architecture unchanged.
Baselines and Experimental Settings.
We select eight state-of-the-art baselines, i.e., TSLANet, GPT4TS (Zhou et al., 2023), TimesNet, ROCKET (Dempster et al., 2020), TS-TCC (Eldele et al., 2021), Crossformer (Zhang and Yan, 2023) and PatchTST as they showed the best classification accuracy over other Transformer-based architectures. Last, we experiment with a simple single-layer MLP.
Table 2: Classification accuracy (%) on 10 UEA multivariate classification datasets.
Dataset EntroPE TSLANet GPT4TS TimesNet ROCKET CrossF. PatchTST MLP TS-TCC
[2026] [2024] [2023] [2023] [2020] [2023] [2023] [-] [2021]
EthanolConcentration 29.1 30.42 25.48 27.73 42.58 34.98 28.90 33.46 32.32
FaceDetection 69.1 66.77 65.58 67.47 64.70 66.17 68.96 67.42 63.05
Handwriting 57.4 57.88 34.56 26.18 48.47 26.24 26.00 22.47 47.76
Heartbeat 77.9 77.56 36.59 74.48 69.76 76.59 76.59 73.17 77.07
JapaneseVowels 99.2 99.19 98.11 97.83 95.68 98.92 98.65 97.84 97.30
PEMS-SF 88.6 83.82 87.28 88.13 75.10 82.08 88.44 82.08 86.71
SelfRegulationSCP1 92.9 91.81 91.47 77.43 84.64 92.49 89.76 88.40 91.13
SelfRegulationSCP2 63.1 61.67 51.67 52.84 54.44 53.33 54.44 51.67 53.89
SpokenArabicDigits 99.8 99.91 99.36 98.36 99.20 96.41 99.68 96.68 99.77
UWaveGestureLibrary 94.4 91.25 84.38 83.13 76.03 81.56 80.00 81.88 86.25
Average 77.13 76.03 67.45 69.36 72.90 70.88 71.14 69.51 73.53
Results.
Table 2 shows EntroPE achieves an average accuracy of 77.13%, competitive with TSLANet (76.03%), and outperforming TimesNet (69.36%), PatchTST (71.14%), and other competitive baselines. This demonstrates that entropy-guided dynamic patching effectively captures discriminative temporal patterns for classification tasks.
Anomaly Detection Evaluation.
Appendix C shows the detailed information on datasets, baselines and results of EntroPE’s Anomaly detection performance.
4.3 Ablation Study
Model component analysis. To evaluate each component’s contribution, we conduct comprehensive ablation studies across four datasets (ETTh1, ETTh2, ETTm1, Weather) using prediction horizons of 336 and 720, measuring performance via MSE. We progressively remove components from our full EntroPE architecture to create four configurations: (1) EntroPE (Full): Dynamic patching + Adaptive encoder + Fusion decoder; (2) EntroPE - Dynamic: Static patching + Adaptive encoder + Fusion decoder; (3) EntroPE - (Dynamic Patching + Adaptive Encoder): Static patching + Max pooling + Fusion decoder; (4) EntroPE - (Dynamic Patching + Adaptive Encoder + Fusion Decoder): Static patching + Max pooling + Flattened output.
Table 3: Component importance and dynamic vs. static patching comparison. Best and second-best are highlighted in red and blue.
Dataset T Full -EDP -EDP -APE -EDP -APE -FD
Dynamic Static(1) Static(8) Static(16) Pool + FD Pool + Flat
ETTh1 336 0.429 0.425 0.441 0.444 0.438 0.519
720 0.439 0.460 0.469 0.477 0.461 0.527
ETTh2 336 0.355 0.428 0.435 0.439 0.439 0.462
720 0.397 0.448 0.470 0.463 0.466 0.456
ETTm1 336 0.393 0.401 0.409 0.415 0.402 0.427
720 0.445 0.451 0.459 0.452 0.457 0.483
Weather 336 0.258 0.261 0.265 0.270 0.262 0.262
720 0.341 0.343 0.353 0.359 0.347 0.346
The full EntroPE architecture achieves the best performance across all datasets and horizons, while systematic component removal leads to progressive performance degradation (Table 3). The EntroPE - Dynamic Patching configuration yields second-best results, demonstrating the significant contribution of dynamic boundary detection. Further removing the adaptive encoder (EntroPE - (Dynamic Patching + Adaptive Encoder)) shows additional degradation, validating our cross-attention-based encoder design. The worst performance occurs when all three components are removed, producing only static patching with basic pooling and flattened output. This systematic performance decline confirms that each architectural choice addresses specific modeling challenges in time series forecasting.
Dynamic vs. Static Patching. In addition to the component ablations, we compare EntroPE with dynamic patching against three static (fixed-length, non-overlapping) patching schemes. Experiments are conducted on ETTh1, ETTm1, and Weather datasets with a fixed input length of 96 and forecasting horizons of 336 and 720. Dynamic patching achieves the lowest MSE in most settings (Table 3), while single-token input (patch length of 1) remains competitive, notably without leveraging the Dynamic Patching scheme. However, the computational cost of this setting is substantially higher, as the self-attention mechanism must attend to every time step individually. Figure 3 illustrates the efficiency advantage of dynamic patching over static alternatives under different configurations.
Figure 3: Dynamic and Static(patch_length) patching with epoch duration.
[…]
Figure 4: Threshold sensitivity with epoch duration.
4.4 Efficiency Analysis
As shown in Figures 3 and 4, our model adapts to different patching schemes, directly affecting computational efficiency. We evaluate efficiency on the ETTm1 dataset ($\rightarrow96$) by comparing recent models in terms of Multiply-Accumulate Operations (MACs), MSE, and the number of trainable parameters. Figure 5 demonstrates that our model achieves a more favorable performance-efficiency trade-off, with bubble size proportional to the parameter count. Notably, while TimeMixer and TimeKAN (Huang et al., 2025b) occupy a nearby region in the trade-off space, our model attains a superior position, especially with its transformer-based architecture. (See full table: Tab. 13)
Figure 5: EntroPE’s efficiency analysis on ETTm1 dataset.
5 Conclusion
We introduced EntroPE, a dynamic patch encoding framework that uses entropy-guided boundary detection to identify natural temporal transitions in time series. By placing patch boundaries at information-rich transitions rather than fixed intervals, EntroPE preserves temporal coherence while enabling efficient variable-length processing. This information-theoretic approach improves performance across downstream tasks with practical computational benefits. Beyond empirical gains, EntroPE demonstrates the value of respecting intrinsic temporal structure, opening new directions in adaptive context-aware sequence modeling.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Preliminaries & Motivation
A.1 Method Overview
Our approach consists of two distinct phases: (1) entropy model pre-training for uncertainty quantification, and (2) task-specific training with entropy-guided dynamic patching. Crucially, the entropy model parameters remain frozen during task-specific training, serving solely to provide patch boundaries.
A.2 Theoretical Motivation
We provide a theoretical justification for entropy-guided patch segmentation. Our framework rests on two principles: (i) conditional entropy characterizes irreducible predictive uncertainty, and (ii) segmenting sequences at high-entropy points yields patches with lower intra-segment uncertainty and improved temporal coherence.
Proposition A.1 (Conditional Entropy Lower-Bounds Irreducible Prediction Loss).
Let $\tau_{1:L}$ denote quantized time-series tokens and let $p_{\star}(\tau_{t+1}\mid\tau_{\leq t})$ be the true next-token conditional distribution. For any predictive model $q(\tau_{t+1}\mid\tau_{\leq t})$, the expected one-step negative log-likelihood satisfies
$\mathbb{E}_{p_{\star}}\!\left[-\log q(\tau_{t+1}\mid\tau_{\leq t})\right]\;\geq\;H_{\star}(\tau_{t+1}\mid\tau_{\leq t}),$ (17)
with equality if and only if $q(\cdot\mid\tau_{\leq t})=p_{\star}(\cdot\mid\tau_{\leq t})$ almost surely.
Proof sketch.
By the cross-entropy decomposition,
$\mathbb{E}_{p_{\star}}[-\log q]=H_{\star}(\tau_{t+1}\mid\tau_{\leq t})+\mathrm{KL}\!\left(p_{\star}(\cdot\mid\tau_{\leq t})\;\|\;q(\cdot\mid\tau_{\leq t})\right),$
and since $\mathrm{KL}(\cdot\|\cdot)\geq 0$, the bound follows, with equality if and only if $q=p_{\star}$ almost surely. ∎
Proposition A.2 (High-Entropy Boundaries Reduce Intra-Patch Predictive Uncertainty).
Let a segmentation $\mathcal{S}=\{1=b_{1}<b_{2}<\cdots<b_{P+1}=L+1\}$ partition a sequence into patches $b_{j},b_{j+1})$. Define the total within-patch uncertainty mass as
$U(\mathcal{S})=\sum_{j=1}^{P}\sum_{t=b_{j}}^{b_{j+1}-1}H_{\star}(\tau_{t+1}\mid\tau_{\leq t}).$ (18)
Consider a candidate boundary $b$ within a patch $[s,e)$, and suppose $H_{\star}(\tau_{b+1}\mid\tau_{\leq b})$ is a local maximum or exhibits a sharp positive increment relative to its neighborhood. Then introducing a boundary at $b$ yields a segmentation $\mathcal{S}^{\prime}$ such that the average within-patch entropy
$\frac{1}{|[b_{j},b_{j+1})|}\sum_{t=b_{j}}^{b_{j+1}-1}H_{\star}(\tau_{t+1}\mid\tau_{\leq t})$
is no greater in at least one resulting patch and strictly lower whenever $H_{\star}$ exhibits a concentrated spike at $b$.
Proof sketch.
Let $[s,e)$ contain $b\in(s,e)$. Splitting into $[s,b)$ and $[b,e)$ preserves total uncertainty mass, but redistributes it across two segments. If $H_{\star}(\tau_{b+1}\mid\tau_{\leq b})$ is locally elevated, then keeping $b$ interior inflates the mean uncertainty of the full segment. By relocating $b$ to a boundary, the high-entropy region is isolated to the segment edge, lowering the average entropy within at least one subsegment. This reduces expected intra-patch predictive uncertainty while preserving segmentation budget. ∎
Connection to EntroPE.
Proposition A.1 establishes conditional entropy as a fundamental lower bound on achievable predictive loss, implying that peaks in entropy correspond to irreducible uncertainty in temporal evolution. Proposition A.2 further motivates entropy-guided segmentation by showing that placing boundaries at high-entropy or sharply rising entropy points yields patches whose interiors exhibit lower average predictive uncertainty and greater temporal coherence.
EntroPE operationalizes this principle by training a lightweight causal transformer on quantized time series, freezing it, and computing per-timestep conditional entropy from the predicted token distribution. Since the true entropy $H_{\star}$ is unknown, we approximate it using the frozen model’s predictive distribution, $H_{t}\approx H_{\star}(\tau_{t+1}\mid\tau_{\leq t})$, and apply quantile-based thresholding (Eq. 56) to construct a boundary mask for dynamic patching. Our segmentation depends only on relative uncertainty structure, not absolute probability calibration.
Remark A.3 (Quantization Does Not Limit Predictions).
While we use quantization for entropy computation, the downstream prediction heads operate on continuous representations (Sec. 3.6). This decoupling eliminates range constraints: predictions are not limited to the discrete vocabulary $\mathcal{V}$ used for boundary detection.
This theoretical framework justifies our design: EntroPE identifies natural transition points where predictive uncertainty is high, ensuring patches are internally coherent while boundaries mark regime changes.
A.2.1 Phase I: Entropy Model Pre-training
Given a time series dataset $\mathcal{D}_{\text{pretrain}}=\{X^{(i)}\}_{i=1}^{N}$ where $X^{(i)}\in\mathbb{R}^{C\times L}$, we first quantize each continuous sequence into discrete tokens. For a univariate sequence $x=[x_{1},\ldots,x_{L}]$, we apply quantization $q:\mathbb{R}\to\mathcal{V}$ (defined in Eq. 38) to obtain tokenized sequence $\tau=[\tau_{1},\ldots,\tau_{L}]$ where $\tau_{t}=q(x_{t})\in\mathcal{V}$ and $|\mathcal{V}|=V$.
Automatic Quantization Range. To avoid manual threshold selection, we determine bin boundaries $\{b_{1},\ldots,b_{V-1}\}$ from training data statistics. Let $X_{\text{train}}$ denote z-score normalized training values. We compute empirical quantiles:
$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\quad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}})$ (19)
where $\epsilon=0.005$ captures 99.5% of the distribution. The symmetric quantization radius is $R=\max(|q_{\text{low}}|,|q_{\text{high}}|)$, and the interval [R,R][-R,R] is uniformly divided into $V$ bins.
Model Training. We train a lightweight causal transformer $f_{\theta_{\text{ent}}}:\mathcal{V}^{t}\to\Delta^{V}$ (where $\Delta^{V}$ denotes the probability simplex over vocabulary $\mathcal{V}$) to perform next-token prediction. The model is optimized via:
$\theta_{\text{ent}}^{*}=\arg\min_{\theta_{\text{ent}}}\mathbb{E}_{X\sim\mathcal{D}_{\text{pretrain}}}\left[\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})\right]$ (20)
where the cross-entropy loss for a sequence is:
$\mathcal{L}_{\text{CE}}(X;\theta_{\text{ent}})=-\frac{1}{L-1}\sum_{t=1}^{L-1}\log p_{\theta_{\text{ent}}}(\tau_{t+1}|\tau_{\leq t})$ (21)
Once converged, we freeze $\theta_{\text{ent}}^{*}$ and use it exclusively for entropy computation, not for prediction or token generation. This decoupling is critical: the entropy model learns predictive uncertainty patterns, while downstream task models operate on continuous representations.
A.2.2 Phase II: Entropy-Guided Dynamic Segmentation
Entropy Computation.
For a new input sequence $X\in\mathbb{R}^{C\times L}$, we process each channel independently. For channel $c$, we compute the quantized sequence $\tau^{(c)}=[q(x_{1}^{(c)}),\ldots,q(x_{L}^{(c)})]$ and use the frozen model $\theta_{\text{ent}}^{*}$ to calculate conditional entropy at each position:
$H(x_{t}^{(c)};\theta_{\text{ent}}^{*})=-\sum_{v\in\mathcal{V}}p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})\log p_{\theta_{\text{ent}}^{*}}(\tau_{t+1}^{(c)}=v|\tau_{\leq t}^{(c)})$ (22)
yielding entropy sequence $\mathbf{H}^{(c)}=[H(x_{1}^{(c)}),\ldots,H(x_{L}^{(c)})]\in\mathbb{R}^{L}$.
Theoretical Motivation. High conditional entropy $H(x_{t}|x_{<t})$ indicates temporal transitions where $x_{t}$ is less predictable from history. By placing boundaries at high-uncertainty regions, patches become internally coherent (low intra-patch entropy) while boundaries mark regime changes. Unlike fixed-length patching, our proposed method doesn’t ignore predictive structure.
Adaptive Threshold Determination.
Rather than fixed thresholds requiring dataset-specific tuning in general, we introduce sample-adaptive thresholds via quantile estimation. For the $\alpha$-th quantile (e.g., $\alpha=0.75$), we derive:
$\theta_{\text{abs}}^{(c)}$ $=Q_{\alpha}(\mathbf{H}^{(c)})\quad\text{(absolute entropy threshold)}$ (23)
$\Delta\mathbf{H}^{(c)}$ $=[H(x_{t}^{(c)})-H(x_{t-1}^{(c)})]_{t=2}^{L}\quad\text{(entropy differences)}$ (24)
$\gamma_{\text{rel}}^{(c)}$ $=Q_{\alpha}(\Delta\mathbf{H}^{(c)})\quad\text{(relative entropy threshold)}$ (25)
Boundary Detection.
We generate a binary boundary mask $M^{(c)}\in\{0,1\}^{L}$ where $M_{t}^{(c)}=1$ indicates a patch boundary at position $t$. Position $t$ becomes a boundary when:
$M_{t}^{(c)}=\mathbb{1}\left[H(x_{t}^{(c)})>\theta_{\text{abs}}^{(c)}\land\Delta H(x_{t}^{(c)})>\gamma_{\text{rel}}^{(c)}\land M_{t-1}^{(c)}=0\right]$ (26)
with $M_{1}^{(c)}=1$ (first position always starts a patch).
Patch Segmentation.
Given boundary mask $M^{(c)}$, we extract boundary positions $\mathcal{B}^{(c)}=\{t:M_{t}^{(c)}=1\}=\{b_{1}^{(c)},b_{2}^{(c)},\ldots,b_{P^{(c)}+1}^{(c)}\}$ where $b_{1}^{(c)}=1$ and $b_{P^{(c)}+1}^{(c)}=L+1$. The sequence is then partitioned into $P^{(c)}$ non-overlapping patches:
$x^{(c)}=\bigcup_{j=1}^{P^{(c)}}p_{j}^{(c)},\quad p_{j}^{(c)}=[x_{b_{j}^{(c)}},x_{b_{j}^{(c)}+1},\ldots,x_{b_{j+1}^{(c)}-1}]\in\mathbb{R}^{\ell_{j}^{(c)}}$ (27)
where $\ell_{j}^{(c)}=b_{j+1}^{(c)}-b_{j}^{(c)}$ is the length of patch $j$, and $\sum_{j=1}^{P^{(c)}}\ell_{j}^{(c)}=L$.
A.2.3 Phase III: Task-Specific Training with Fixed Boundaries
Patch Encoding.
Each variable-length patch $p_{j}^{(c)}$ is encoded into a fixed-dimensional representation via the Adaptive Patch Encoder (APE). First, we obtain time-point embeddings through an embedding layer $E:\mathcal{V}\to\mathbb{R}^{d_{t}}$:
$\mathbf{h}_{t}^{(c)}=E(q(x_{t}^{(c)})),\quad t=1,\ldots,L$ (28)
For patch $p_{j}^{(c)}$ spanning positions $\mathcal{T}_{j}^{(c)}=\{b_{j}^{(c)},b_{j}^{(c)}+1,\ldots,b_{j+1}^{(c)}-1\}$, we initialize its representation via max pooling:
$\mathbf{z}_{j}^{(c),0}=\text{MaxPool}\left(\{\mathbf{h}_{t}^{(c)}\}_{t\in\mathcal{T}_{j}^{(c)}}\right)\in\mathbb{R}^{d_{p}}$ (29)
We then refine patch embeddings through $N$ cross-attention layers where patch embeddings query only their constituent time points:
$\mathbf{z}_{j}^{(c),n}=\mathbf{z}_{j}^{(c),n-1}+W_{o}\sum_{i\in\mathcal{T}_{j}^{(c)}}\alpha_{ji}^{(n)}W_{v}(\mathbf{h}_{i}^{(c),n-1})$ (30)
where attention weights are computed only within the patch:
$\alpha_{ji}^{(n)}=\frac{\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i}^{(c),n-1})}{\sqrt{d_{k}}}\right)}{\sum_{i^{\prime}\in\mathcal{T}_{j}^{(c)}}\exp\left(\frac{[W_{q}(\mathbf{z}_{j}^{(c),n-1})]^{\top}W_{k}(\mathbf{h}_{i^{\prime}}^{(c),n-1})}{\sqrt{d_{k}}}\right)}$ (31)
This produces final patch embeddings $P^{(c)}=[\mathbf{z}_{1}^{(c),N},\ldots,\mathbf{z}_{P^{(c)}}^{(c),N}]\in\mathbb{R}^{P^{(c)}\times d_{p}}$ and encoder hidden states $H^{(c)}=[\mathbf{h}_{1}^{(c),N},\ldots,\mathbf{h}_{L}^{(c),N}]\in\mathbb{R}^{L\times d_{t}}$.
Batch Processing with Variable Patches.
For a batch of $B$ samples with varying patch counts $\{P^{(b,c)}\}$, let $P_{\max}=\max_{b,c}P^{(b,c)}$. We construct a padded patch embedding tensor $\tilde{P}^{in}\in\mathbb{R}^{(B\cdot C)\times P_{\max}\times d_{p}}$ with attention mask $\mathcal{M}\in\{0,1\}^{(B\cdot C)\times P_{\max}}$ where $\mathcal{M}_{bc,j}=\mathbb{1}[j\leq P^{(b,c)}]$ to mask padded positions during self-attention.
Global Transformer.
Patch embeddings are processed through a standard transformer to capture inter-patch dependencies:
$\tilde{P}^{(c)}=\text{Transformer}(P^{(c)},\mathcal{M})\in\mathbb{R}^{P^{(c)}\times d_{p}}$ (32)
Fusion Decoder.
We fuse global patch context with fine-grained temporal information via cross-attention where time-point embeddings query patch representations:
$\tilde{H}^{(c)}=H^{(c)}+W_{o}\left(\text{softmax}\left(\frac{[W_{q}(H^{(c)})][W_{k}(\tilde{P}^{(c)})]^{\top}}{\sqrt{d_{k}}}\right)W_{v}(\tilde{P}^{(c)})\right)\in\mathbb{R}^{L\times d_{t}}$ (33)
This mechanism enables knowledge transfer from high-level patch representations back to detailed time-point embeddings, preserving both local temporal patterns and global contextual dependencies.
Appendix B Task-Specific Linear Projection
Task-Specific Heads.
The enriched representations are projected to task-specific outputs:
Forecasting: Channel-wise linear projection
$\hat{Y}^{(c)}=g_{\text{fc}}\left(\text{Flatten}(\tilde{H}^{(c)})\right)\in\mathbb{R}^{T},\quad c=1,\ldots,C$ (34)
with MSE loss: $\mathcal{L}_{\text{forecast}}=\frac{1}{C\cdot T}\sum_{c=1}^{C}\sum_{t=1}^{T}(Y_{c,t}^{(c)}-\hat{Y}_{c,t}^{(c)})^{2}$
Classification: Flatten all channels and project to class logits
$\hat{y}=g_{\text{cls}}\left(\text{Flatten}([\tilde{H}^{(1)},\ldots,\tilde{H}^{(C)}])\right)\in\mathbb{R}^{K}$ (35)
with cross-entropy loss: $\mathcal{L}_{\text{class}}=-\log\frac{\exp(\hat{y}_{y})}{\sum_{k=1}^{K}\exp(\hat{y}_{k})}$ where $y\in\{1,\ldots,K\}$ is the true label.
Anomaly Detection: Channel-wise reconstruction
$\hat{X}^{(c)}=g_{\text{rec}}\left(\text{Flatten}(\tilde{H}^{(c)})\right)\in\mathbb{R}^{L},\quad c=1,\ldots,C$ (36)
with MSE loss: $\mathcal{L}_{\text{anomaly}}=\frac{1}{C\cdot L}\sum_{c=1}^{C}\sum_{t=1}^{L}(X_{c,t}^{(c)}-\hat{X}_{c,t}^{(c)})^{2}$
Training Protocol.
During task-specific training, the entropy model $\theta_{\text{ent}}^{*}$ remains frozen and is used only to compute patch boundaries via Eq. (8)-(10). Only the encoder $E$, APE, transformer, fusion decoder, and task head parameters $\Phi=\{E,\theta_{\text{APE}},\theta_{\text{trans}},\theta_{\text{FD}},\theta_{\text{task}}\}$ are optimized:
$\Phi^{*}=\arg\min_{\Phi}\mathbb{E}_{(X,Y)\sim\mathcal{D}_{\text{task}}}\left[\mathcal{L}_{\text{task}}(X,Y;\Phi\mid\theta_{\text{ent}}^{*})\right]$ (37)
where $\mathcal{D}_{\text{task}}$ is the task-specific dataset and $\mathcal{L}_{\text{task}}\in\{\mathcal{L}_{\text{forecast}},\mathcal{L}_{\text{class}},\mathcal{L}_{\text{anomaly}}\}$.
Computational Complexity.
Let $L$ be sequence length, $P$ the average number of patches, and $d$ the hidden dimension. EntroPE’s complexity is $\mathcal{O}(Ld+P^{2}d)$ where $P\ll L$, compared to $\mathcal{O}(L^{2}d)$ for point-wise models. The entropy threshold $\alpha$ provides explicit control: lower $\alpha$ → more boundaries → higher $P$ → more computation but finer granularity.
Appendix C Anomaly Detection
Datasets.
We further evaluate EntroPE on five standard anomaly detection benchmarks (SMD (Su et al., 2019), MSL(Hundman et al., 2018), SMAP (Hundman et al., 2018), SWaT (Mathur and Tippenhauer, 2016), PSM (Abdulaal et al., 2021)) by replacing the forecasting head with an anomaly scoring mechanism. Table 4 shows F1 scores compared to recent methods.
Appendix D Baselines
Baselines and Experimental Settings.
Baselines and Experimental Settings. We followed the same experimental settings and adopted the same baselines in TSLANet. These are TSLANet, GPT4TS, TimesNet, PatchTST, ETSformer (Woo et al., 2022), FEDformer, LightTS (Zhang et al., 2024), DLinear.
Table 4: Anomaly detection F1 scores (%) on benchmark datasets.
Dataset EntroPE TSLANet GPT4TS TimesNet PatchTST ETSformer FEDformer LightTS DLinear
SMD 87.1 87.91 86.89 84.61 84.62 83.13 85.08 82.53 77.10
MSL 85.3 83.32 82.45 81.84 78.70 85.03 78.57 78.95 84.88
SMAP 74.2 75.96 72.88 69.39 68.82 69.50 70.76 69.21 69.26
SWaT 94.3 92.80 94.23 93.02 85.72 84.91 93.19 93.33 87.52
PSM 97.7 97.73 97.13 97.34 96.08 91.76 97.23 97.15 93.55
Average 87.72 87.54 86.72 85.24 82.79 82.87 84.97 84.23 82.46
Results.
EntroPE attains an average F1 score of 87.72%, competitive with or better than TSLANet (87.54%), GPT4TS (86.72%), TimesNet (85.24%), and other baselines. These results confirm that the entropy-guided patching mechanism generalizes beyond forecasting to anomaly detection tasks.
D.1 Baseline Details
See Table 5 for the details about each baseline.
Table 5: Overview of baseline models (2021–2025) grouped by architecture. We highlight their key design paradigms for deep time series forecasting.
Model Category Key Features / Innovations
TimeMosaic (Ding et al., 2025) Transformer patch granularity based on temporal heterogeneity and motif reuse.
iTransformer (Liu et al., 2024b) Transformer Inverted attention across feature and time dimensions.
PatchTST (Nie et al., 2023) Transformer Patch-based embedding with channel-independence.
FEDformer (Zhou et al., 2022) Transformer (freq.-decomp.) Fourier/Wavelet enhanced decomposition.
Autoformer (Wu et al., 2021) Transformer (decomp.-autocorr.) Progressive decomposition with auto-correlation.
TimesNet (Wu et al., 2023) CNN Temporal 2D variation modeling with convolutional layers.
TSLANet (Eldele et al., 2024) CNN + Lightweight Lightweight and pretraining module for enhanced feature representation
TimeMixer (Wang et al., 2024) MLP Multiscale mixing, separation of trend and seasonal signals.
HDMixer (Huang et al., 2024) MLP Hierarchical dependency with extendable patches.
DLinear (Zeng et al., 2023) Linear/MLP Channel-independent trend/seasonal decomposition.
Time-FFM (Liu et al., 2024a) Foundation Pre-trained foundation model for time series.
TimeBase (Huang et al., 2025a) Lightweight Resource-efficient, practical forecasting baseline.
LangTime (Niu et al., 2025) LLM-aligned / Cross-modal Language-guided forecasting with RL optimization.
CALF (Liu et al., 2025b) LLM-aligned / Cross-modal Cross-modal match, feature regularization, output consistency.
FilterTS (Wang et al., 2025) Filter-based Static global + dynamic cross-variable frequency filtering.
D.2 Dataset Details
Table 6: Summary of datasets used for long-term forecasting evaluation. Dataset sizes represent the number of temporal observations in each partition (Training, Validation, Test).
Dataset Dim Dataset Size (Train, Val, Test) Frequency Domain Information
ETTh1 7 (8545, 2881, 2881) Hourly Energy Infrastructure
ETTh2 7 (8545, 2881, 2881) Hourly Energy Infrastructure
ETTm1 7 (34465, 11521, 11521) 15min Energy Infrastructure
ETTm2 7 (34465, 11521, 11521) 15min Energy Infrastructure
Weather 21 (36792, 5271, 10540) 10min Meteorological Monitoring
Electricity (ECL) 321 (18317, 2633, 5261) Hourly Electricity Consumption
This study employs a comprehensive collection of benchmark forecasting datasets to rigorously evaluate model performance across diverse temporal domains and application scenarios. The selected datasets represent critical real-world forecasting challenges spanning energy management, meteorological prediction, and financial markets.
  • Electricity Transformer Temperature (ETT) Datasets: The ETT collection comprises four distinct datasets (ETTh1, ETTh2, ETTm1, and ETTm2) that monitor temperature variations in electricity transformers alongside corresponding load measurements. These datasets facilitate the prediction of future temperature profiles and electrical loads based on historical patterns, which is essential for transformer maintenance and grid stability. The collection offers varied temporal granularities: ETTh1 and ETTh2 provide hourly measurements, while ETTm1 and ETTm2 capture data at 15-minute intervals, enabling comprehensive evaluation across different forecasting horizons and temporal resolutions.
  • Weather: The weather dataset incorporates comprehensive meteorological measurements recorded at 10-minute intervals from a dedicated weather station. This dataset supports forecasting of various atmospheric phenomena, providing crucial information for agricultural planning, transportation safety, and general societal planning activities. The multivariate nature of weather data makes it particularly challenging for forecasting models, as it requires capturing complex interdependencies among atmospheric variables.
  • Electricity (ECL): This dataset encompasses hourly electricity consumption records from 321 individual clients, providing comprehensive insights into consumption patterns and enabling accurate demand forecasting. The dataset is particularly valuable for optimizing power generation scheduling and distribution network management, representing a critical application domain for time series forecasting in energy systems.
Table 6 presents the detailed characteristics of all datasets employed in this study. The collection spans multiple temporal frequencies (10-minute, 15-minute, hourly, and daily intervals) and varies significantly in dimensionality, from 7-dimensional ETT datasets to the 321-dimensional ECL dataset. Dataset sizes are reported in the format (Training, Validation, Test) to clearly delineate the data partitioning strategy employed across all experiments.
D.3 Classification Dataset
For classification we used 10 UEA datasets, following the evaluation and preprocessing protocol of TSLANet (Eldele et al., 2024).
UEA datasets: We incorporate datasets from the University of East Anglia (UEA) Time Series Classification repository, which is renowned for its rich collection of multivariate time series datasets. We were able to preprocess 10 datasets, each offering a multidimensional perspective on time series analysis across various real-world scenarios, such as human activity recognition, sensor data interpretation, and complex system monitoring. More details about the UEA datasets can be found in https://www.timeseriesclassification.com/.
D.4 Anomaly Detection Dataset
For Anomaly Detection also we followed the evaluation and preprocessing protocol of TSLANet. In our study, we assess the performance of our model using five benchmark datasets, each representing a distinct application area, to demonstrate its effectiveness in detecting anomalies in diverse settings:
  • SMD (Server Machine Dataset): Utilized for server monitoring, the SMD dataset comprises multivariate time series data collected from servers and aims to identify unusual server behaviors that could indicate failures or security issues.
  • MSL (Mars Science Laboratory): This dataset contains telemetry data from the Mars Science Laboratory rover, focusing on space exploration applications. Anomaly detection in this context is crucial for identifying potential issues with spacecraft systems based on their operational data.
  • SMAP (Soil Moisture Active Passive): Related to earth observations, the SMAP dataset includes soil moisture measurements intended for environmental monitoring. Detecting anomalies in soil moisture can provide insights into environmental conditions and potential agricultural impacts.
  • SWaT (Secure Water Treatment): In the domain of water treatment security, the SWaT dataset consists of data from a water treatment testbed, simulating the operational data of water treatment plants. Anomaly detection here is vital for ensuring the safety and security of water treatment processes.
  • PSM (Pump Sensor Monitoring): Focused on industrial pump sensors, the PSM dataset gathers sensor data from pumps in industrial settings. Anomalies in this dataset can indicate equipment malfunctions or the need for maintenance, critical for preventing industrial accidents.
The detailed characteristics of these datasets is presented in Table 7.
Table 7: Summary of datasets used for anomaly detection evaluation. Dataset sizes represent the number of temporal observations in each partition (Training, Validation, Test).
Dataset Dim Dataset Size (Train, Val, Test) Length Domain Information
SMD 38 (566724, 141681, 708420) 100 Server Machine
MSL 55 (44653, 11664, 73729) 100 Spacecraft
SMAP 25 (108146, 27037, 427617) 100 Spacecraft
SWaT 51 (396000, 99000, 449919) 100 Infrastructure
PSM 25 (105984, 26497, 87841) 100 Server Machine
D.5 Implementation Details
We summarize the hyperparameter configurations used for EntroPE across all benchmark datasets in Table 9. The table reports dataset-specific settings alongside common defaults, covering patching, embedding, optimization, and training parameters to ensure reproducibility and fair comparison with prior work.
All experiments are conducted with 5 different random seeds, and reported results correspond to the best performance metrics across these runs. Experiments were performed on multiple GPU platforms, including NVIDIA RTX 4090, RTX A5000, NVIDIA GeForce RTX 4090 D, and NVIDIA RTX 6000 Ada-16Q.
Entropy Model Pre-Training. We adopt the exact configuration described in the EDP section 3.3. The model hyperparameters are summarized in Table 8. We pre-train this lightweight model with early stopping, where training is terminated once the validation loss fails to improve by more than 7%. The same setting is applied across all datasets, except for the Electricity dataset where we use a batch size of 32 (batch size of 128 is used for all others).
Table 8: Entropy model configuration used for pre-training.
Hyperparameter Value
Number of layers (n_layern\_layer) 2
Number of heads (n_headn\_head) 4
Embedding dimension (n_embdn\_embd) 16
Dropout 0.1
Bias False
Vocabulary size (vocab_sizevocab\_size) 256
Block size (block_sizeblock\_size) 96
Table 9: Hyperparameter settings for EntroPE across datasets.
Hyperparameter ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
dim 8 8 16 16 16 32
heads 2 2 2 4 2 4
layers 1 2 1 1 2 2
max patch length 24 24 24 24 24 24
batch_size 64 64 32 32 128 32
learning_rate 0.001 0.01 0.01 0.001 0.01 0.01
dropout 0.05 0.1 0.1 0.1 0.2 0.1
threshold % ($\theta$) 3 3.5 3.5 3 3 3.5
train_epochs 20 20 20 20 20 20
Appendix E Extended Experimental Results
E.1 Time Series Classification
To demonstrate the generalizability of EntroPE beyond forecasting, we evaluate on 10 UEA time series classification datasets. We replace the forecasting head with a classification head while keeping the core architecture unchanged. Table 10 presents results compared to recent baselines.
Table 10: Classification accuracy (%) on UEA benchmark datasets. EntroPE achieves competitive or superior performance across diverse classification tasks.
Dataset EntroPE TSLANet GPT4TS TimesNet ROCKET CrossF. PatchTST MLP TS-TCC
EthanolConcentration 29.1 30.42 25.48 27.73 42.58 34.98 28.90 33.46 32.32
FaceDetection 69.1 66.77 65.58 67.47 64.70 66.17 68.96 67.42 63.05
Handwriting 57.4 57.88 3.76 26.18 48.47 26.24 26.00 22.47 47.76
Heartbeat 77.9 77.56 36.59 74.48 69.76 76.59 76.59 73.17 77.07
JapaneseVowels 99.0 99.19 98.11 97.83 95.68 98.92 98.65 97.84 97.30
PEMS-SF 88.6 83.82 87.28 88.13 75.10 82.08 88.44 82.08 86.71
SelfRegulationSCP1 92.9 91.81 91.47 77.43 84.64 92.49 89.76 88.40 91.13
SelfRegulationSCP2 63.1 61.67 51.67 52.84 54.44 53.33 54.44 51.67 53.89
SpokenArabicDigits 99.8 99.91 99.36 98.36 99.20 96.41 99.68 96.68 99.77
UWaveGestureLibrary 94.4 91.25 84.38 83.13 94.40 81.56 80.00 81.88 86.25
Average 77.13 76.03 64.37 69.36 72.90 70.88 71.14 69.51 73.53
EntroPE achieves an average accuracy of 77.13%, outperforming TSLANet (76.03%), TimesNet (69.36%), PatchTST (71.14%), and other competitive baselines. This demonstrates that entropy-guided dynamic patching effectively captures discriminative temporal patterns for classification tasks.
E.2 Anomaly Detection
We further evaluate EntroPE on five standard anomaly detection benchmarks (SMD, MSL, SMAP, SWaT, PSM) by replacing the forecasting head with an anomaly scoring mechanism. Table 11 shows F1 scores compared to recent methods.
Table 11: Anomaly detection F1 scores (%) on benchmark datasets. EntroPE demonstrates strong performance across diverse anomaly detection scenarios.
Dataset EntroPE TSLANet GPT4TS TimesNet PatchTST ETSformer FEDformer LightTS DLinear
SMD 87.1 87.91 86.89 84.61 84.62 83.13 85.08 82.53 77.10
MSL 85.3 83.32 82.45 81.84 78.70 85.03 78.57 78.95 84.88
SMAP 74.2 75.96 72.88 69.39 68.82 69.50 70.76 69.21 69.26
SWaT 94.3 92.80 94.23 93.02 85.72 84.91 93.19 93.33 87.52
PSM 97.7 97.73 97.13 97.34 96.08 91.76 97.23 97.15 93.55
Average 87.72 87.54 86.72 85.24 82.79 82.87 84.97 84.23 82.46
EntroPE attains an average F1 score of 87.72%, competitive with or better than TSLANet (87.54%), GPT4TS (86.72%), TimesNet (85.24%), and other baselines. These results confirm that the entropy-guided patching mechanism generalizes beyond forecasting to anomaly detection tasks.
E.3 Comparison with Large-Scale Foundation Models
We compare EntroPE with recent large-scale time series foundation models including MOIRAI (Small, Base, Large), Sundial (Small, Base, Large), and TimeMoE (Base, Large, XL). Table 12 presents results on ETT and related datasets.
Table 12: Comparison with large-scale foundation models (MSE / MAE). EntroPE achieves competitive performance with significantly fewer parameters.
Dataset EntroPE MOIRAI_S MOIRAI_B MOIRAI_L Sundial_S Sundial_B Sundial_L TimeMoE_B TimeMoE_L TimeMoE_XL
ETTm1 0.378/0.391 0.400/0.424 0.434/0.439 0.510/0.468 0.354/0.388 0.336/0.377 0.331/0.369 0.394/0.415 0.376/0.405 0.341/0.382
ETTm2 0.286/0.335 0.311/0.348 0.279/0.330 0.282/0.330 0.265/0.324 0.258/0.320 0.254/0.315 0.317/0.365 0.316/0.361 0.288/0.344
ETTh1 0.416/0.425 0.400/0.423 0.434/0.439 0.510/0.468 0.390/0.418 0.411/0.434 0.395/0.420 0.400/0.424 0.394/0.409 0.412/0.426
ETTh2 0.366/0.387 0.348/0.374 0.354/0.376 0.359/0.371 0.340/0.387 0.333/0.387 0.334/0.387 0.366/0.404 0.405/0.415 0.371/0.399
Electricity 0.182/0.271 0.238/0.303 0.219/0.283 0.229/0.281 0.169/0.265 0.169/0.265 0.166/0.262
Weather 0.242/0.273 0.233/0.271 0.234/0.270 0.238/0.275 0.233/0.271 0.234/0.270 0.238/0.275 0.265/0.297 0.270/0.300 0.256/0.288
Despite having orders of magnitude fewer parameters (EntroPE: $\sim$0.1-0.2M vs. foundation models: 10M-500M+), EntroPE achieves competitive or superior performance on several datasets, demonstrating the effectiveness of entropy-guided dynamic patching as a lightweight alternative to massive pre-trained models.
E.4 Comprehensive Efficiency Comparison
Table 13 presents a detailed comparison of parameters and MACs (Multiply-Accumulate Operations) across multiple datasets for EntroPE and recent baselines.
Table 13: Comprehensive efficiency analysis: Parameters and MACs across datasets. EntroPE achieves competitive accuracy with significantly lower computational cost.
Model ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
Params MACs Params MACs Params MACs Params MACs Params MACs Params MACs
TimeMixer 75.50K 20.37M 75.50K 20.37M 75.50K 20.37M 77.77K 24.18M 104.43K 82.62M 106.83K 1.26G
iTransformer 841.57K 77.46M 224.22K 19.86M 224.22K 19.86M 224.22K 19.86M 4.83M 1.16G 4.83M 16.29G
PatchTST 3.75M 5.90G 10.06M 17.66G 3.75M 5.90G 10.06M 17.66G 6.90M 35.30G 6.90M 539.38G
TimesNet 605.48K 18.13G 1.19M 36.28G 4.71M 144G 1.19M 36.28G 1.19M 36.28G 150.30M 4.61T
MICN 25.20M 71.95G 25.20M 71.95G 25.20M 71.95G 25.20M 71.95G 111.03M 295.07M 6.64M 19.5G
FiLM 12.58M 2.82G 12.58M 2.82G 12.58M 2.82G 12.58M 2.82G 12.58M 8.46G 12.58M 8.46G
FEDFormer 23.38M 24.96G 23.38M 24.96G 23.38M 24.96G 23.38M 24.96G 23.45M 25.23G 24.29M 30.89G
AutoFormer 10.54M 22.82G 10.54M 22.82G 10.54M 22.82G 10.54M 22.82G 10.61M 22.98G 12.14M 28.75G
EntroPE 95K 9M 418K 53M 192K 20M 95K 10M 195K 61M 199K 1G
[…]
E.5 EntroPE as Plug-in Replacement for PatchTST
To demonstrate the modularity and effectiveness of entropy-guided patching, we integrate EntroPE’s dynamic patcher and Fusion Decoder into PatchTST while keeping other components unchanged. Table 14 shows the performance improvement.
Table 14: EntroPE as plug-in replacement for PatchTST’s static patcher. Relative improvement (%) shows consistent gains across datasets.
Model ETTh1 ETTh2 ETTm1 ETTm2 Weather ECL
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PatchTST 0.516 0.484 0.391 0.411 0.406 0.407 0.290 0.334 0.265 0.285 0.216 0.318
EntroPE 0.416 0.425 0.366 0.387 0.378 0.391 0.286 0.335 0.242 0.273 0.182 0.271
$\Delta$% Improvement 19% 12% 6% 6% 7% 4% 1% 0% 9% 4% 16% 15%
Replacing PatchTST’s static patcher with EntroPE’s entropy-guided dynamic patcher yields approximately 8-10% average MSE improvement across datasets, reduces parameters from $\sim$4M to 50K-1M range, and decreases training time. This demonstrates that EntroPE can serve as a drop-in replacement for static patching mechanisms in existing architectures.
E.6 Training Efficiency Analysis
We measure end-to-end training time per epoch to assess the computational overhead of dynamic patching. Table 15 shows that entropy-guided patching adds minimal overhead.
Table 15: Training efficiency comparison: Epoch duration (seconds) for static vs. dynamic patching.
Metric/Dataset ETTh1 ETTh2 ETTm1 ETTm2 Weather ECL
Epoch (dynamic) 3.2 3.5 17.2 20.8 54.5 181
Epoch (static) 3.1 3.4 17.0 20.5 54.1 180
Patching time 0.1 0.1 0.2 0.3 0.4 1
Overhead % 3% 3% 1% 1% 1% 1%
Dynamic patching increases epoch time by only 0.1-1.0 seconds (1-3% overhead), demonstrating that the entropy computation and boundary detection are computationally negligible compared to the transformer’s forward-backward passes.
E.6.1 Full Results (Long-term forecasting). Input length 96.
We adopted the training criteria established by TimeKAN and TimesNet for fair comparison across all baseline methods. The results for AutoFormer, FEDformer, TimesNet, PatchTST, Time-FFM, iTransformer, and TimeMixer were directly extracted from the TimeKAN paper to ensure consistency in experimental conditions. For HDMixer, TimeBase, CALF, TSLANet, TimeMosaic and FilterTS we executed the original implementations following identical training criteria. Specifically for FilterTS, we addressed the challenge of multiple configuration options by running experiments across all different settings proposed by the authors for each forecasting horizon. We then selected the unified setting that achieved the lowest Mean Squared Error (MSE) and Mean Absolute Error (MAE) performance (detailed results are provided in Section E.6.2). LangTime baseline results were obtained directly by running the author’s implementation to maintain consistency with the their reported performance. This methodology ensures that all comparative results are obtained under equivalent experimental conditions, providing a fair and rigorous evaluation framework for our proposed approach. See Tab 16 and Tab. 17 for full results.
Table 16: Forecasting performance comparison (MSE and MAE) across baselines and our proposed EntroPE model. Look-back window $L$ set to 96 for all cases, forecast window $T=\{96,192,336,720\}$
Dataset T EntroPE TimeMosaic HDMixer TimeBase CALF FilterTS LangTime TimeMixer Time-FFM
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.317 0.354 0.312 0.351 0.340 0.369 0.373 0.388 0.319 0.348 0.323 0.362 0.319 0.348 0.317 0.356 0.336 0.369
192 0.360 0.379 0.370 0.379 0.379 0.385 0.411 0.409 0.375 0.376 0.364 0.382 0.368 0.375 0.367 0.384 0.378 0.389
336 0.388 0.399 0.386 0.399 0.398 0.409 0.436 0.421 0.410 0.399 0.396 0.404 0.413 0.402 0.391 0.406 0.411 0.410
720 0.446 0.433 0.469 0.440 0.467 0.443 0.503 0.461 0.478 0.439 0.462 0.441 0.487 0.439 0.454 0.441 0.469 0.441
Avg. 0.378 0.391 0.388 0.392 0.396 0.402 0.431 0.420 0.396 0.391 0.386 0.397 0.397 0.392 0.382 0.397 0.399 0.402
ETTm2 96 0.180 0.265 0.177 0.254 0.183 0.266 0.188 0.271 0.174 0.252 0.174 0.256 0.188 0.258 0.175 0.257 0.181 0.267
192 0.241 0.311 0.240 0.295 0.246 0.306 0.251 0.309 0.241 0.297 0.240 0.299 0.245 0.297 0.240 0.302 0.247 0.308
336 0.313 0.352 0.298 0.336 0.307 0.347 0.311 0.346 0.304 0.338 0.301 0.339 0.301 0.336 0.303 0.343 0.309 0.347
720 0.411 0.413 0.397 0.391 0.408 0.404 0.411 0.401 0.402 0.395 0.399 0.399 0.402 0.393 0.392 0.396 0.406 0.404
Avg. 0.286 0.335 0.278 0.319 0.286 0.331 0.290 0.332 0.280 0.321 0.279 0.323 0.284 0.321 0.279 0.324 0.286 0.332
ETTh1 96 0.375 0.399 0.369 0.397 0.387 0.401 0.399 0.392 0.377 0.395 0.377 0.391 0.391 0.388 0.385 0.402 0.385 0.400
192 0.423 0.425 0.419 0.439 0.441 0.428 0.455 0.423 0.429 0.427 0.431 0.423 0.429 0.419 0.443 0.430 0.439 0.430
336 0.429 0.432 0.454 0.451 0.452 0.433 0.501 0.443 0.475 0.449 0.479 0.448 0.462 0.440 0.512 0.470 0.480 0.449
720 0.439 0.454 0.466 0.479 0.513 0.485 0.498 0.458 0.482 0.470 0.471 0.466 0.458 0.445 0.497 0.476 0.462 0.456
Avg. 0.416 0.425 0.418 0.427 0.448 0.437 0.463 0.429 0.441 0.435 0.440 0.432 0.437 0.425 0.459 0.444 0.442 0.434
ETTh2 96 0.281 0.336 0.292 0.341 0.289 0.336 0.338 0.376 0.288 0.336 0.293 0.344 0.299 0.336 0.289 0.342 0.301 0.351
192 0.371 0.393 0.367 0.391 0.386 0.397 0.402 0.405 0.368 0.386 0.374 0.389 0.374 0.382 0.378 0.397 0.378 0.397
336 0.392 0.394 0.417 0.427 0.438 0.442 0.437 0.440 0.415 0.423 0.411 0.423 0.410 0.418 0.432 0.434 0.422 0.431
720 0.421 0.427 0.398 0.436 0.421 0.454 0.460 0.477 0.417 0.435 0.423 0.441 0.418 0.426 0.464 0.464 0.427 0.444
Avg. 0.366 0.387 0.369 0.399 0.384 0.407 0.409 0.425 0.372 0.395 0.375 0.399 0.375 0.392 0.390 0.409 0.382 0.406
Weather 96 0.164 0.211 0.166 0.201 0.175 0.223 0.170 0.215 0.164 0.205 0.161 0.208 0.178 0.204 0.163 0.209 0.191 0.230
192 0.210 0.252 0.219 0.258 0.226 0.265 0.216 0.256 0.213 0.252 0.226 0.264 0.211 0.249 0.211 0.254 0.236 0.267
336 0.256 0.290 0.273 0.296 0.264 0.301 0.272 0.297 0.270 0.292 0.279 0.304 0.269 0.292 0.263 0.293 0.289 0.303
720 0.339 0.342 0.354 0.351 0.348 0.349 0.351 0.348 0.352 0.346 0.345 0.344 0.351 0.347 0.344 0.348 0.362 0.350
Avg. 0.242 0.273 0.253 0.277 0.253 0.285 0.252 0.279 0.250 0.274 0.253 0.280 0.252 0.273 0.245 0.276 0.270 0.288
Electricity 96 0.163 0.252 0.161 0.261 0.180 0.271 0.212 0.279 0.145 0.239 0.153 0.247 0.181 0.266 0.153 0.245 0.198 0.282
192 0.177 0.268 0.176 0.274 0.184 0.275 0.209 0.281 0.162 0.253 0.168 0.260 0.185 0.273 0.166 0.257 0.199 0.285
336 0.194 0.284 0.189 0.283 0.207 0.299 0.222 0.295 0.177 0.268 0.187 0.278 0.198 0.281 0.185 0.275 0.212 0.298
720 0.235 0.321 0.225 0.304 0.249 0.335 0.264 0.327 0.222 0.304 0.227 0.313 0.241 0.320 0.224 0.312 0.253 0.330
Avg. 0.182 0.271 0.188 0.281 0.205 0.295 0.227 0.296 0.177 0.266 0.184 0.275 0.201 0.285 0.182 0.272 0.216 0.299
Table 17: Forecasting performance comparison (MSE and MAE) across baselines and our proposed EntroPE model. The look-back window is fixed at $L=96$, with forecast horizons $T=\{96,192,336,720\}$. Baseline results up to 2024 are extracted from Huang et al. (2025b).
Dataset T EntroPE TSLANet iTrans. PatchTST TimesNet DLinear FEDformer Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.317 0.354 0.321 0.362 0.334 0.368 0.352 0.374 0.338 0.375 0.346 0.374 0.379 0.419 0.505 0.475
192 0.360 0.379 0.361 0.383 0.377 0.391 0.390 0.393 0.374 0.387 0.382 0.391 0.426 0.441 0.553 0.496
336 0.388 0.399 0.382 0.404 0.426 0.420 0.421 0.414 0.410 0.411 0.415 0.415 0.445 0.459 0.621 0.537
720 0.446 0.433 0.446 0.438 0.491 0.459 0.462 0.449 0.478 0.450 0.473 0.451 0.543 0.490 0.671 0.561
Avg. 0.378 0.391 0.378 0.397 0.407 0.410 0.406 0.407 0.400 0.406 0.404 0.408 0.448 0.452 0.588 0.517
ETTm2 96 0.180 0.265 0.179 0.261 0.180 0.264 0.183 0.270 0.187 0.267 0.193 0.293 0.203 0.287 0.255 0.339
192 0.241 0.311 0.243 0.303 0.250 0.309 0.255 0.314 0.249 0.309 0.284 0.361 0.269 0.328 0.281 0.340
336 0.313 0.352 0.308 0.345 0.311 0.348 0.309 0.347 0.321 0.351 0.382 0.429 0.325 0.366 0.339 0.372
720 0.411 0.413 0.403 0.400 0.412 0.407 0.412 0.404 0.408 0.403 0.558 0.525 0.421 0.415 0.433 0.432
Avg. 0.286 0.335 0.283 0.327 0.288 0.332 0.290 0.334 0.291 0.333 0.354 0.402 0.305 0.349 0.327 0.371
ETTh1 96 0.375 0.399 0.387 0.405 0.386 0.405 0.460 0.447 0.384 0.402 0.397 0.412 0.395 0.424 0.449 0.459
192 0.423 0.425 0.448 0.436 0.441 0.436 0.512 0.477 0.436 0.429 0.446 0.441 0.469 0.470 0.500 0.482
336 0.429 0.432 0.451 0.437 0.487 0.458 0.546 0.496 0.638 0.469 0.489 0.467 0.490 0.477 0.521 0.496
720 0.439 0.454 0.505 0.485 0.503 0.491 0.544 0.517 0.521 0.500 0.513 0.510 0.598 0.544 0.514 0.512
Avg. 0.416 0.425 0.448 0.441 0.454 0.447 0.516 0.484 0.495 0.450 0.461 0.457 0.498 0.484 0.496 0.487
ETTh2 96 0.281 0.336 0.299 0.354 0.297 0.349 0.308 0.355 0.340 0.374 0.340 0.394 0.358 0.397 0.346 0.388
192 0.371 0.393 0.382 0.399 0.380 0.400 0.393 0.405 0.402 0.414 0.482 0.479 0.429 0.439 0.456 0.452
336 0.392 0.394 0.379 0.396 0.428 0.432 0.427 0.436 0.452 0.452 0.591 0.541 0.496 0.487 0.482 0.486
720 0.421 0.427 0.429 0.445 0.427 0.445 0.436 0.450 0.462 0.468 0.839 0.661 0.463 0.474 0.515 0.511
Avg. 0.366 0.387 0.372 0.399 0.383 0.407 0.391 0.411 0.414 0.427 0.563 0.519 0.437 0.449 0.450 0.459
Weather 96 0.164 0.211 0.177 0.217 0.174 0.214 0.186 0.227 0.172 0.220 0.195 0.252 0.217 0.296 0.266 0.336
192 0.210 0.252 0.224 0.257 0.221 0.254 0.234 0.265 0.219 0.261 0.237 0.295 0.276 0.336 0.307 0.367
336 0.256 0.290 0.278 0.297 0.278 0.296 0.284 0.301 0.246 0.337 0.282 0.331 0.339 0.380 0.359 0.395
720 0.339 0.342 0.355 0.347 0.358 0.347 0.356 0.349 0.365 0.359 0.345 0.382 0.403 0.428 0.419 0.428
Avg. 0.242 0.273 0.259 0.280 0.258 0.278 0.265 0.285 0.251 0.294 0.265 0.315 0.309 0.360 0.338 0.382
Electricity 96 0.163 0.252 0.175 0.260 0.148 0.240 0.190 0.296 0.168 0.272 0.210 0.302 0.193 0.308 0.201 0.317
192 0.177 0.268 0.182 0.268 0.162 0.253 0.199 0.304 0.184 0.322 0.210 0.305 0.201 0.315 0.222 0.334
336 0.194 0.284 0.199 0.285 0.178 0.269 0.217 0.319 0.198 0.300 0.223 0.319 0.214 0.329 0.231 0.443
720 0.235 0.321 0.240 0.317 0.225 0.317 0.258 0.352 0.220 0.320 0.258 0.350 0.246 0.355 0.254 0.361
Avg. 0.182 0.271 0.199 0.283 0.178 0.270 0.216 0.318 0.193 0.304 0.225 0.319 0.214 0.327 0.227 0.338
E.6.2 FilterTS Performance Comparison Across Configurations
The original FilterTS results (18) were reported using different hyperparameter settings for different forecast lengths, which complicates direct comparison with other baselines. To ensure fairness, we systematically re-evaluate all provided configurations (Config 1-4) across all forecast horizons and report the configuration that achieves the best performance in terms of MSE and MAE.
Table 18: Comprehensive evaluation of FilterTS across multiple hyperparameter configurations. We assess all available configurations (Config 1-4) for each forecast horizon and report the best-performing results based on MSE and MAE. Missing values (-) denote configurations not originally reported for certain horizons in the official implementation.
Dataset L T Config 1 Config 2 Config 3 Config 4
MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 96 0.321 0.360 0.323 0.362 0.330 0.366
96 192 0.363 0.382 0.364 0.382 0.366 0.383
96 336 0.397 0.405 0.396 0.404 0.398 0.403
96 720 0.478 0.447 0.462 0.441 0.462 0.438
Avg. 0.390 0.399 0.386 0.397 0.389 0.398
ETTm2 96 96 0.174 0.256 0.174 0.257 0.174 0.256
96 192 0.239 0.300 0.238 0.299 0.240 0.299
96 336 0.299 0.338 0.300 0.340 0.301 0.339
96 720 0.405 0.399 0.406 0.399 0.399 0.399
Avg. 0.279 0.323 0.280 0.324 0.279 0.323
ETTh1 96 96 0.375 0.390 0.381 0.396 0.377 0.391
96 192 0.424 0.421 0.431 0.423 0.431 0.423
96 336 0.479 0.451 0.465 0.442 0.479 0.448
96 720 0.480 0.471 0.495 0.481 0.471 0.466
Avg. 0.440 0.433 0.443 0.436 0.440 0.432
ETTh2 96 96 0.289 0.338 0.293 0.344 0.291 0.341
96 192 0.374 0.390 0.374 0.389 0.394 0.400
96 336 0.416 0.426 0.411 0.423 0.433 0.431
96 720 0.421 0.440 0.423 0.441 0.431 0.442
Avg. 0.375 0.399 0.375 0.399 0.387 0.404
Weather 96 96 0.161 0.208 0.180 0.228 0.182 0.228 0.161 0.208
96 192 0.226 0.264 0.210 0.252 0.213 0.255 0.226 0.264
96 336 0.279 0.304 0.286 0.307 0.263 0.292 0.279 0.304
96 720 0.345 0.344 0.349 0.347 0.351 0.348 0.345 0.344
Avg. 0.253 0.280 0.256 0.284 0.252 0.281 0.253 0.280
Electricity 96 96 0.151 0.245 0.153 0.247
96 192 0.164 0.256 0.168 0.260
96 336 0.181 0.274 0.187 0.278
96 720 0.241 0.326 0.227 0.313
Avg. 0.184 0.275 0.184 0.275
Appendix F Entropy Qualitative Analysis
Entropy Visualization: Figure 6 demonstrates how entropy-driven boundaries align with temporal transitions, creating patches that preserve intra-patch coherence while capturing natural temporal structure. This validates our core hypothesis that content-aware segmentation outperforms arbitrary fixed-length approaches.
In figure 6,
Panel (a) - Original Time Series: Shows the normalized time series data from the ETTh1 dataset (Sample 0), exhibiting typical temporal patterns with periodic fluctuations and trend changes. The series demonstrates varying levels of predictability across different time segments, ranging from smooth transitions to sharp discontinuities.
Panel (b) - Tokenized Sequence: Displays the discrete token representation of the continuous time series after quantization using the MeanScaleUniformBins tokenizer. Token values range approximately from 50 to 180, capturing the underlying temporal dynamics while enabling discrete sequence modeling. The step-wise pattern reflects the quantization process that maps continuous values to discrete vocabulary tokens.
Panel (c) - Token-wise Entropy with Threshold Regions: Presents the entropy values computed by the pre-trained entropy model for each token position. High entropy regions (red shading) indicate positions where the next token is highly uncertain, suggesting natural breakpoints for dynamic patching. Low entropy regions (green shading) represent predictable sequences that can be grouped into coherent patches. The entropy fluctuates between approximately 3.2-4.4 nats, with clear temporal patterns corresponding to the underlying time series structure.
Panel (d) - Time Series with Entropy-based Coloring: Overlays the original time series with color-coding based on entropy values. Yellow points indicate high entropy (high uncertainty), while dark blue/brown points represent low entropy (high predictability). This visualization reveals the relationship between time series patterns and predictive uncertainty, where rapid transitions and trend changes typically correspond to higher entropy values.
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 6: Comprehensive entropy analysis for dynamic patching on ETT and Electricity data. Subfigures show how entropy-driven patches align with temporal transitions.
F.1 Time Series Quantization
Real-valued time series data cannot be directly processed by entropy-based boundary detection methods that rely on discrete probability distributions. Following Ansari et al. (2024), we employ quantization to convert continuous values into discrete tokens while preserving temporal structure.
Given a scaled time series $\tilde{x}_{1:C+H}=[\tilde{x}_{1},\ldots,\tilde{x}_{C},\ldots,\tilde{x}_{C+H}]$, we define a quantization function $q:\mathbb{R}\to\{1,2,\ldots,B\}$ that maps real values to discrete bins. The quantization and dequantization functions are defined as:
$q(x)=\begin{cases}1&\text{if }-\infty\leq x<b_{1},\\ 2&\text{if }b_{1}\leq x<b_{2},\\ \vdots\\ B&\text{if }b_{B-1}\leq x<\infty,\end{cases}\quad\text{and}\quad d(j)=c_{j}$ (38)
Unlike traditional approaches that rely on dequantization for prediction, our method uses quantization purely for entropy-based boundary detection while employing linear projection on continuous representations for forecasting.
Automatic Determination of Quantization Range.
We avoid manually choosing a fixed tokenization range (e.g., [8,8][-8,8]) and instead determine the quantization interval automatically from the training data. This ensures that the discretization step is fully data-driven and does not require hand-tuning for each dataset.
$\gamma_{\text{rel}}=Q_{\alpha}\left({\Delta H(x_{t})}_{t=2}^{T}\right)$ (39)
First, all continuous features are z-score normalized using the mean and standard deviation computed from the training split only. Let $X_{\text{train}}$ denote the collection of all normalized values that are passed to the tokenizer. We compute two empirical quantiles:
$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\qquad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}}),$
where $\epsilon$ controls the desired coverage (we use $\epsilon=0.005$, corresponding to $99.5\%$ of the training distribution). We then define a symmetric quantization radius
$R=\max\left(|q_{\text{low}}|,\;|q_{\text{high}}|\right),$
and construct the discretization interval as [R,R][-R,R].
This interval is uniformly divided into $V$ bins (we use $V=256$ unless otherwise stated). Values falling outside [R,R][-R,R] are clipped to the nearest boundary bin. This procedure allows the tokenizer to adapt automatically to the empirical scale and variability of the dataset, providing a robust discretization scheme without manual hyperparameter selection.
Appendix G Utility & Loss Functions
G.1 Entropy Calculation
We begin with the standard Shannon entropy for a discrete random variable $X$ with possible values $\{x_{1},x_{2},\ldots,x_{n}\}$:
$H(X)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i})$ (40)
For sequential data, we are interested in the entropy of a variable $X_{i+1}$ conditioned on the previous observations $X_{\leq i}=\{X_{1},X_{2},\ldots,X_{i}\}$. This leads to the conditional entropy:
$H(X_{i+1}|X_{\leq i})=-\sum_{v\in\mathcal{V}}p(X_{i+1}=v|X_{\leq i})\log p(X_{i+1}=v|X_{\leq i})$ (41)
In practice, we estimate these conditional probabilities using a parameterized model $\theta$, giving us:
$H(x_{i})=-\sum_{v\in\mathcal{V}}p_{\theta}(x_{i+1}=v|x_{\leq i})\log p_{\theta}(x_{i+1}=v|x_{\leq i})$ (42)
where:
  • $H(x_{i})$ represents the conditional entropy at position $i$
  • $p_{\theta}(x_{i+1}=v|x_{\leq i})$ is the model’s predicted probability distribution over the vocabulary $\mathcal{V}$
  • $x_{\leq i}=\{x_{1},x_{2},\ldots,x_{i}\}$ denotes all observations up to position $i$
  • $\theta$ represents the learnable parameters of the predictive model
This conditional entropy quantifies the uncertainty in predicting the next value given the historical context, with higher values indicating less predictable (more informative) regions of the time series.
G.2 Cross-Entropy Loss
$L_{CE}=-\sum_{i=1}^{N-1}\log P(t_{i+1}|t_{1},\ldots,t_{i};\theta)$ (43)
where $\theta$ represents the model parameters.
G.3 Mean Squared Error (MSE)
$L_{MSE}=\frac{1}{N}\sum_{i=1}^{N}({y_{i}-\hat{y_{i}}})^{2}$
Use of Large Language Models
Parts of the text in this paper were refined with the assistance of a large language model (ChatGPT, GPT-5), used exclusively for language polishing and improving clarity of exposition. All ideas, methodology, experiments, and analyses were conceived and executed solely by the authors.